# First-passage properties of the jump process with a drift. The general case

### v1

@@ 1. Introduction
Imagine a bakery with a single counter where a cashier serves customers who arrive at random times, each requiring a varying amount of time to complete their order. The cashier works at a consistent pace, attending to one customer at a time in the order they arrive, with no overtaking allowed. This setup, that we have so frivolously described, actually represents a fundamental model in queuing theory. Specifically, if both the inter-arrival times and the amount of time it takes to serve a customer are independent identically distributed random variables, this process is usually referred to as a G/G/1 queue (see [1] for a pedagogical introduction).
A simple observation is that if many customers arrive with complex orders, the queue will grow indefinitely. In contrast, if few customers arrive or orders are simple to fulfill, the cashier will remain idle most of the time. So it is natural to wonder whether the queue depletes and, if so, how much time does it take? This is essentially the question we address in the present paper.
@@
To formulate the problem in a more formal manner, we consider a jump process with a drift, i.e., a stochastic process $X(t)$ evolving on the real line through a combination of deterministic negative linear drift and random positive jumps (see Fig. 1). In the <<bakery>> picture, $X(t)$ is the total amount of requests piled in the queue, the linear decay segments correspond to the service of customers, while the jumps represent the sporadic arrivals of newcomers. The dynamics of $X(t)$ unfold as follows:
@@
    The process starts at an initial position $X_{0}$, and immediately experiences a jump of random amplitude $M_{1}$. It then drifts toward the origin at constant speed $\alpha$ for a random time interval $t_{1}$, after which the next jump $M_{2}$ occurs.
@@
In [36] we proposed an alternative approach that does not rely on the renewal equation. Specifically, we constructed a mapping of the original continuous-time process onto an effective discrete-time random walk. This mapping allowed us to obtain a closed-form representation of the Laplace transform of the joint probability distribution of $\tau$ and $n$, by applying the Pollaczek-Spitzer formula [37, 38, 39], a powerful result in discrete random walk theory. More precisely, we used a generalization [40] to the asymmetric random walk of the formula obtained by Ivanov in [41]. This approach was then successfully applied to two exactly solvable cases with Poisson arrivals, where explicit solutions could be obtained and compared against standard renewal equation techniques.
The present work tackles a significantly more challenging general case, where explicit solutions are no longer available. Here, we consider a rather general class of distributions requiring that $p(t)$ and $q(M)$ be light-tailed with smooth probability densities, thereby moving far beyond the assumption of Poisson arrivals. In this general case, the closed-form integral representations of the relevant Laplace transforms are still valid, but extracting the asymptotic behavior now requires analytic continuation techniques and singularity analysis. In this work we provide such analysis and derive several exact asymptotic results for this general setting.
The main results concern the behavior of the survival probabilities, i.e., the probability that the process stays positive up to time $\tau$ (or up to the $n$-th jump). We show that there are three distinct regimes determined by the drift strength. In the survival regime (weak drift), the process has a finite probability of never crossing the origin, and survival probabilities decay exponentially to a positive constant. In the absorption regime (strong drift), the first passage event occurs with certainty, and survival probabilities decay exponentially to zero. We compute these exponential decay rates explicitly for arbitrary light-tailed distributions with smooth densities. At the critical point separating these regimes, exponential decay is replaced by algebraic decay. Such behavior mirrors that obtained in [36] for two exactly solvable cases.
@@
The paper is organized as follows: In Section 2 we define the model more formally and recall the representation for the triple Laplace transform (with respect to $\tau$, $n$ and $X_{0}$) of the joint probability distribution of $\tau$ and $n$ obtained in [36]. In Section 3 we study the properties of this transform in the $\lambda$-plane ($\lambda$ being the Laplace dual of $X_{0}$) and perform analytic continuation. Sections 4, 5 and 6 analyze the asymptotic behavior of the survival probabilities: at infinite time, up to the $n$-th jump and up to time $\tau$, respectively. Section 7 concerns the asymptotic expansions of the means and variances of $\tau$ and $n$ in the absorption regime. Finally, we conclude in Section 8.
@@ 2. The model and the main results
In this paper we consider the process (1) in a rather general setup. That is, we do not specify the distributions $p(t)$ and $q(M)$ and instead only require them to have two properties:
@@
    All moments of both $M$ and $t$ are finite (both distributions are light-tailed).
@@
    The densities $p(t)$ and $q(M)$ are continuously differentiable functions, supported either on a finite interval $0,\ell]$ or on the positive half-line $0,\infty)$.
@@
The transition between these regimes can be understood by examining the average displacement per cycle. Each cycle consists of a jump of size $M$ followed by drift of $\alpha t$ toward the origin. The sign of the average displacement determines the regime:
@@
We treat the survival probability at infinite time as an <<order parameter>> separating the survival regime, where $S_{\infty}(X_{0})>0$, from the absorption regime, where $S_{\infty}(X_{0})=0$, with $\alpha$ being the <<control parameter>>. The case in which $\alpha=\alpha_{c}$ we refer to as the critical point (see Fig. 2 for a schematic representation).
@@
The function $F(k;\rho)$ entering (9) is a Fourier transform of the probability distribution of the steps in the effective random walk. In terms of the original probability distributions $p(t)$ and $q(M)$ it reads
@@
The joint probability distribution $\mathbb{P}[\tau,n\,|\,X_{0}]$ could, in principle, be found by inverting the transform in (7). This is, however, a very thorny path and there is little reason to hope that it can be traversed for arbitrary distributions. Instead we study the analytic structure of the triple Laplace transform $\hat{Q}(\rho,s\,|\,\lambda)$ and obtain the asymptotic results. The main caveat is that the parameters $\lambda$, $\rho$ and $s$ originate from the Laplace transforms; hence the representation (9) is valid for $\lambda>0$, $\rho>0$, and $s\in(0,1)$. However, extracting the asymptotic behavior of the probabilities requires extending this representation to larger domains through proper analytic continuation.
@@
Finally, we make two remarks. First, we note that alternative representations for survival probabilities of discrete-time random walks have recently been obtained using Riemann-Hilbert techniques [57], providing a complementary approach to the Pollaczek-Spitzer formula. Second, one of the two exactly solvable cases in [36] takes $q(M)$ to be a delta function; hence, the smoothness requirement is not fulfilled, but the detailed computation using (8) can still be performed. The reason is that this requirement is mostly technical and can be lifted entirely for a significant portion of the derivations. In addition, the results can readily be extended to cases where only the first three moments of $M$ and $t$ are finite. Keeping this level of precision, however, comes at the cost of significantly more technical detail without offering much additional insight. For this reason, we choose to sacrifice a small degree of generality in favor of clarity, and throughout the paper we always assume that both distributions $p(t)$ and $q(M)$ have smooth densities and that all their moments are finite.
@@ 2.1. The main results
When the drift is weak ($\alpha<\alpha_{c}$), there is a finite probability that the process remains positive indefinitely, $S_{\infty}(X_{0})>0$. The key result here is explicit expressions for the exponential decay rates that govern the approach to this asymptotic value.
@@
$\frac{1}{\xi_{n}(\alpha)}=-\log\left[\min_{k_{I}\in(\zeta_{-},\zeta_{+})}F(\mathrm{i}k_{I};0)\right],\qquad\frac{1}{\xi_{\tau}(\alpha)}=-\rho_{*},$ (14)
where $\rho_{*}<0$ is determined by the condition
@@
The minimization is performed over purely imaginary arguments within the region where $\mathrm{Im}(k)\in(\zeta_{-},\zeta_{+})$}, which is the maximal strip in the complex $k$-plane where the Fourier transform $F(k;\rho)$ defined in (11) is analytic (see Section 3.1 for details).
Additionally, we obtain the asymptotic behavior of $S_{\infty}(X_{0})$ itself. If the process starts close to the origin, then
$S_{\infty}(X_{0})\underset{X_{0}\to 0}{\sim}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\langle(M-\alpha t)^{2}\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\log\left|1-F(k;0)\right|\right).$ (16)
For initial positions far from the origin, we obtain exponential behavior
@@
Similarly, when $\tau\to\infty$ and $X_{0}\to\infty$ with $X_{0}/\sqrt{\tau}$ fixed:
@@
The rates given by the same expressions as in the survival regime (14).
In this regime the moments of $\tau$ and $n$ are finite, and their Laplace transform with respect to $X_{0}$ can be found from (8). Here we obtain the first two terms in the asymptotic expansions as $X_{0}\to\infty$ and $X_{0}\to 0$ for the mean and variance of both $n$ and $\tau$:
@@
Let us conclude the presentation of the results by providing some motivation for constructing expansions beyond the leading order. While subleading coefficients might appear as technical details, in some cases they carry significant physical meaning. A notable example is the constant $U_{0}$ in the critical regime expansion (110), which in the symmetric random walk case is closely related to the Milne extrapolation length in radiative transport theory [58] and trapping problem for Rayleigh flights [59] (see [51] for a more comprehensive overview). Similar constants also find applications in packing problems in computer science [60, 61]. Since these corrections are nontrivial to obtain, we believe it is worthwhile to present here a systematic derivation based on Mellin transform techniques, which provides a robust framework for extracting such asymptotic expansions.

### v2

@@ 1. Introduction
Imagine a bakery with a single counter where a cashier serves customers who arrive at random times, each requiring a varying amount of time to complete their order. The cashier works at a consistent pace, attending to one customer at a time in the order they arrive, with no overtaking allowed. This setup, that we have so frivolously described, actually represents a fundamental model in queuing theory. Specifically, if both the inter-arrival times and the amount of time it takes to serve a customer are independent and identically distributed random variables, this process is usually referred to as a G/G/1 queue (see [undef] for a pedagogical introduction).
A simple observation is that if many customers arrive with complex orders, the queue will grow indefinitely. In contrast, if few customers arrive or orders are simple to fulfill, the cashier will remain idle most of the time. So it is natural to wonder whether the queue depletes and, if so, how much time it takes? This is essentially the question we address in the present paper.
@@
To formulate the problem in a more formal manner, we consider a jump process with a drift, i.e., a stochastic process $X(t)$ evolving on the real line through a combination of deterministic negative linear drift and random positive jumps (see Fig. 1). In the <<bakery>> picture, $X(t)$ is the total amount of requests piled in the queue, the linear decay segments correspond to the service of customers, while the jumps represent the sporadic arrivals of new customers. The dynamics of $X(t)$ unfold as follows:
@@
    The process starts at an initial position $X_{0}$, and immediately experiences a jump of random amplitude $M_{1}$. It then drifts toward the origin with the constant velocity $\alpha$ for a random time interval $t_{1}$, after which the next jump $M_{2}$ occurs.
@@
[…]
In [undefal] we proposed an alternative approach that does not rely on the renewal equation. Specifically, we constructed a mapping of the original continuous-time process onto an effective discrete-time random walk. This mapping allowed us to obtain a closed-form representation of the Laplace transform of the joint probability distribution of $\tau$ and $n$, by applying the Pollaczek-Spitzer formula [undefam, undefan, undefao], a powerful result in discrete random walk theory. More precisely, we used a generalization of Ivanov’s formula [undefap] to the asymmetric random walk [undefaq]. This approach was then successfully applied to two exactly solvable cases with Poisson arrivals, where explicit solutions could be obtained and compared against standard renewal equation techniques.
The present work tackles a significantly more challenging general case, where explicit solutions are no longer available. Here, we consider a rather general class of distributions, namely light-tailed $p(t)$ and $q(M)$ with smooth probability densities, thereby moving far beyond the assumption of Poisson arrivals. In this general case, the closed-form integral representations of the relevant Laplace transforms are still valid, but extracting the asymptotic behavior now requires analytic continuation techniques and singularity analysis. In this work, we provide such analysis and derive several exact asymptotic results for this general setting.
The main results concern the behavior of the survival probabilities, i.e., the probability that the process stays positive up to time $\tau$ (or up to the $n$-th jump). We show that there are three distinct regimes determined by the drift strength. In the survival regime (weak drift), the process has a finite probability of never crossing the origin, and survival probabilities decay exponentially to a positive constant. In the absorption regime (strong drift), the first passage event occurs with certainty, and survival probabilities decay exponentially to zero. We compute these exponential decay rates explicitly for arbitrary light-tailed distributions with smooth densities. At the critical point separating these regimes, exponential decay is replaced by algebraic decay.
@@
As a continuation of [undefal], this paper yields results of similar structure, making the differences worth emphasizing explicitly. The key novelty is that we treat general inter-arrival and jump distributions where the defining integrals cannot be evaluated explicitly. Consequently, the asymptotic analysis is significantly more involved and requires fundamentally different methodology. Regarding the similarity of the results, this is itself a central contribution of this work: we demonstrate that the regime trichotomy (survival, absorption, critical) holds across all light-tailed distributions with smooth densities. While regime boundaries and decay rates depend on $p(t)$ and $q(M)$, the qualitative structure is universal, revealing the two exactly solvable cases to be exemplars of general behavior rather than artifacts of Poisson arrivals.
The paper is organized as follows: In Section 2, we formally define the model and recall from [undefal] the representation for the triple Laplace transform (with respect to $\tau$, $n$, and $X_{0}$) of the joint probability distribution of the first-passage observables $\tau$ and $n$. Then we briefly summarize the main results derived in this paper. Section 3 presents a detailed analysis of a toy example through which we illustrate the analytic continuation techniques used throughout the paper. In Section 4, we study the properties of the Laplace transform in the $\lambda$-plane ($\lambda$ being the Laplace dual of $X_{0}$) and perform analytic continuation. Sections 5, 6, and 7 analyze the asymptotic behavior of the survival probabilities: at infinite time, up to the $n$-th jump, and up to time $\tau$, respectively. Section 8 concerns the asymptotic expansions of the means and the variances of $\tau$ and $n$ in the absorption regime. Finally, we conclude in Section 9.
@@ 2. The model and the main results
In this paper we consider the process (1) in a rather general setup. That is, we do not specify the distributions $p(t)$ and $q(M)$ and instead only require them to satisfy two conditions:
@@
    All moments of the distributions $p(t)$ and $q(M)$ are finite (both distributions are light-tailed).
@@
    The densities $p(t)$ and $q(M)$ are continuously differentiable functions, supported on either a finite interval $0,\ell]$ or the positive half-line $0,\infty)$.
@@
The transition between these regimes can be understood by examining the average displacement per cycle. Each cycle consists of a jump of size $M$ followed by drift of magnitude $\alpha t$ toward the origin. The sign of the average displacement determines the regime:
@@
We treat the survival probability at infinite time as an <<order parameter>> separating the survival regime, where $S_{\infty}(X_{0})>0$, from the absorption regime, where $S_{\infty}(X_{0})=0$, with $\alpha$ being the <<control parameter>>. The case $\alpha=\alpha_{c}$ we refer to as the critical point (see Fig. 2 for a schematic representation).
@@
The function $F(k;\rho)$ appearing in (9) is a Fourier transform of the probability distribution of the steps in the effective random walk. In terms of the original probability distributions $p(t)$ and $q(M)$ it reads
@@
The joint probability distribution $\mathbb{P}[\tau,n\,|\,X_{0}]$ could, in principle, be found by inverting the transform in (7). This is, however, a very thorny path and there is little reason to hope that it can be traversed for arbitrary distributions. Instead, we study the analytic structure of the triple Laplace transform $\hat{Q}(\rho,s\,|\,\lambda)$ and obtain asymptotic results. The main caveat is that the parameters $\lambda$, $\rho$, and $s$ originate from the Laplace transforms; hence the representation (9) is valid for $\lambda>0$, $\rho>0$, and $s\in(0,1)$. However, extracting the asymptotic behavior of the probabilities requires extending this representation to larger domains through proper analytic continuation.
@@
Finally, two remarks are in order. First, alternative representations for survival probabilities of discrete-time random walks have recently been obtained using Riemann-Hilbert techniques [undefaag], providing a complementary approach to the Pollaczek-Spitzer formula. Second, one of the two exactly solvable cases in [undefal] takes $q(M)$ to be a delta function; hence, the smoothness requirement is not fulfilled, but the detailed computation using (8) can still be performed. The reason is that this requirement is mostly technical and can be lifted entirely for a significant portion of the derivations. In addition, the results can readily be extended to cases where only the first three moments of $M$ and $t$ are finite. However, tracking exactly which assumptions are needed for each result would require a substantial amount of bookkeeping without offering much additional insight. For this reason, we choose to sacrifice a small degree of generality in favor of clarity, and throughout the paper we always assume that both distributions $p(t)$ and $q(M)$ have smooth densities and that all their moments are finite.
@@ 2.1. The main results
When the drift is weak ($\alpha<\alpha_{c}$), there is a finite probability that the process remains positive indefinitely, $S_{\infty}(X_{0})>0$. The key result here are explicit expressions for the exponential decay rates that govern the approach to this asymptotic value.
@@
$\frac{1}{\xi_{n}(\alpha)}=\log s_{*},\qquad\frac{1}{\xi_{\tau}(\alpha)}=-\rho_{*},$ (14)
where
$s_{*}=\left[\min_{k_{I}\in(\zeta_{-},\zeta_{+})}F(\mathrm{i}k_{I};0)\right]^{-1},$ (15)
and $\rho_{*}<0$ is determined by the condition
@@
The minimization is performed over purely imaginary arguments within the region where $\mathrm{Im}(k)\in(\zeta_{-},\zeta_{+})$}, which is the maximal strip in the complex $k$-plane where the Fourier transform $F(k;\rho)$, defined in (11), is analytic (see Section 4.1 for details).
Additionally, we obtain the asymptotic behavior of $S_{\infty}(X_{0})$ itself. If the process starts close to the origin, we obtain the expansion
$S_{\infty}(X_{0})\underset{X_{0}\to 0}{\sim}\varsigma_{0}+\varsigma_{1}X_{0},$ (17)
where the coefficients $\varsigma_{0,1}$ are given in (75). For initial positions far from the origin, we instead find exponential behavior
@@
Similarly, when $\tau\to\infty$ and $X_{0}\to\infty$ with $X_{0}/\sqrt{\tau}$ fixed, we have:
@@
The rates are given by the same expressions as in the survival regime (14).
In this regime, the moments of $\tau$ and $n$ are finite, and their Laplace transform with respect to $X_{0}$ can be found from (8). Here we obtain the first two terms in the asymptotic expansions as $X_{0}\to\infty$ and $X_{0}\to 0$ for the mean and variance of both $n$ and $\tau$:
@@
Let us conclude the presentation of the results by providing some motivation for constructing expansions beyond the leading order. While subleading coefficients might appear to be technical details, in some cases they carry significant physical meaning. A notable example is the constant $U_{0}$ in the critical regime expansion (119), which in the symmetric random walk case is closely related to the Milne extrapolation length in radiative transport theory [undefaah] and the trapping problem for Rayleigh flights [undefaai] (see [undefaaa] for a more comprehensive overview). Similar constants also find applications in packing problems in computer science [undefaaj, undefaak]. Since these corrections are nontrivial to obtain, we believe it is worthwhile to present here a systematic derivation based on Mellin transform techniques, which provides a robust framework for extracting such asymptotic expansions.
3. A toy example
Before analyzing $\phi^{\pm}(\lambda;\rho,s)$, it is instructive to consider a simplified toy problem through which to introduce the techniques we shall use and provide the reader with the intuition behind them. Specifically, we define two auxiliary functions
$\psi^{\pm}(\lambda;A,B)=\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\,\frac{1}{\lambda\pm\mathrm{i}k}\log\big[(k-\mathrm{i}A)(k+\mathrm{i}B)\big]\right],$ (29)
where $A$ and $B$ are positive real parameters and the integral is understood as a principal value. In the complex $k$-plane, the analytic structure of the integrands entering (29) is apparent: they have a pole at $k=\mp\mathrm{i}\lambda$ and two logarithmic cuts going from $\mathrm{i}A$ to $\mathrm{i}\infty$ and from $\mathrm{i}B$ to $\mathrm{i}\infty$. In Section 4, we demonstrate that $\phi^{\pm}(\lambda;\rho,s)$ each has a pole and two branch cuts — though the branch points have more intricate expressions. In subsequent sections, we show that these are exactly the singularities responsible for the asymptotic behavior.
Since the functions $\psi^{\pm}(\lambda;A,B)$ resemble $\phi^{\pm}(\lambda;\rho,s)$ in structure, but are sufficiently simple to allow explicit evaluation for all complex $\lambda$, they offer an illustrative example exposing the difference between a naive extension of (29) and a proper analytic continuation. Moreover, there is a pragmatic motive for this detour: similar integrals will reappear in the analysis of first-passage properties, and it is prudent to compute them here, killing two birds with one well-aimed stone.
Recall that the full analysis of $\phi^{\pm}(\lambda;\rho,s)$ requires continuation not only in the $\lambda$-plane, but also in the $s$- and $\rho$-planes. To illustrate these additional continuations, we now introduce $s$-dependence by hand. Specifically, rather than treating $A$ and $B$ as fixed constants, we parametrize them as functions of an auxiliary variable $s$:
$A\equiv A(s)=k_{*}+\sqrt{s_{*}-s},\quad B\equiv B(s)=-k_{*}+\sqrt{s_{*}-s},\quad s\in(0,s_{*}-k_{*}^{2}),$ (30)
where $k_{*}>0$ and $s_{*}>1$ are parameters satisfying $k_{*}^{2}<s_{*}$. At this point such parametrization may appear arbitrary, but it will be justified in Section 4. We choose to denote the parameter by $s$, since, as demonstrated in Section 6, it plays the same role as the Laplace conjugate of $n$ in (7).
Below we first evaluate $\psi^{\pm}(\lambda;A,B)$ explicitly for all complex $\lambda$ and fixed positive values of $A$ and $B$. Once the exact representations are found, performing analytic continuation in the $s$- and $\lambda$-planes is straightforward. In particular, we shall see that the analytic continuations indeed differ from the naive extensions of (29). Our goal here, however, is not to construct such continuations from the explicit results, but rather to perform the continuation directly at the level of the integral representation (29).
In short, the continuation in the $\lambda$-plane requires tracking the movement of the pole at $k=\mp\mathrm{i}\lambda$, and the continuation in the $s$-plane necessitates analyzing the movement of the branch points at $\mathrm{i}A(s)$ and $\mathrm{i}B(s)$ accompanied by appropriate deformations of the integration contour. This is the same procedure used in Section 4 and Section 6 for the continuation of the original functions $\phi^{\pm}(\lambda;\rho,s)$ in the $\lambda$- and $s$-planes, respectively. Regarding the continuation in the $\rho$-plane, the procedure is essentially the same as in the $s$-plane, and we thus do not illustrate it with the toy example.
Explicit evaluation
We begin by evaluating $\psi^{\pm}(\lambda;A,B)$ for all complex $\lambda$ with fixed positive $A$ and $B$. The standard method for evaluating the integrals of the form (29) involves extending them into the complex $k$-plane and closing the contour either in the lower or in the upper half-plane. If the contribution from the semicircle vanishes at infinity, the integral is determined by the singularities inside the contour. However, in the form (29) it is not apparent why the contributions from the arcs decay sufficiently fast (faster than $\left|k\right|^{-1}$). For this reason, we modify this representation so that the standard procedure applies.
$\psi^{\pm}(\lambda)=\exp\left[-\frac{1}{4\pi}\int_{-\infty}^{\infty}\mathrm{d}k\left\{\frac{\log\left[(k-\mathrm{i}A)(k+\mathrm{i}B)\right]}{\lambda\pm\mathrm{i}k}+\frac{\log\left[(k+\mathrm{i}A)(k-\mathrm{i}B)\right]}{\lambda\mp\mathrm{i}k}\right\}\right],$ (31)
where, to simplify notation, we temporarily suppress the explicit dependence on the parameters $A$ and $B$. Rearranging the terms in the integrand we obtain
Applying (34) to the integrals in (33) after some algebraic manipulations, we find that
[…]
From (35) it is evident that the analytic continuations $\psi_{\text{a.c.}}^{\pm}(\lambda)$ of the functions $\psi^{\pm}(\lambda)$ from the positive half of the real line $\lambda\in(0,\infty)$ to the full $\lambda$-plane are given by:
Not only do the functions $\psi^{\pm}(\lambda)$, computed in (35) via the original integral expression (29) for arbitrary $\lambda$, differ from their analytic continuations (36), they are in fact discontinuous at $\mathrm{Re}(\lambda)=0$. With this explicit computation in hand, we now show how to perform the analytic continuation at the level of the integral expression (29).
Continuation in the $\lambda$-plane
We first focus on the function $\psi^{+}(\lambda)$. The analytic structure of the integrand for $\psi^{+}(\lambda)$ in the $k$-plane is relatively simple: a branch cut from the logarithm and a pole at $k=\mathrm{i}\lambda$. When $\mathrm{Re}(\lambda)\to 0^{+}$, this pole approaches the real axis from above, falling on the integration contour when $\mathrm{Re}(\lambda)=0$ and crossing it to the lower half-plane for $\mathrm{Re}(\lambda)<0$. This is exactly the mechanism responsible for the discontinuity of $\psi^{+}(\lambda)$.
[…]
An important remark here is that the above procedure implicitly assumes that the deformed contour $\mathcal{C}$ does not cross the branch cut of the logarithm. This assumption is, however, easily lifted. Consider a strip $\mathrm{Im}(k)\in(-a,a)$, where $a=\min(A,B)$. Within this strip, the logarithmic term in the integrand (29) is analytic. If $\left|\,\mathrm{Re}(\lambda)\right|<a$, then the deformed contour lies within $\mathrm{Im}(k)\in(-a,a)$ where there are no singularities. Hence the analytic continuation (39) is valid in this strip. To extend the continuation to the full region $\mathrm{Re}(\lambda)<0$, we combine (39) with (29) to obtain the functional relation
[…]
Clearly (41) and (39) coincide.
Continuation in the $s$-plane
We now demonstrate how to perform analytic continuation in the $s$-plane if the parameters $A$ and $B$ are not fixed but are instead $s$-dependent and given by (30). From (36), the analytic continuation is straightforward,
$\psi^{\pm}_{\text{a.c.}}(\lambda;s)\equiv\psi^{\pm}_{\text{a.c.}}(\lambda;A(s),B(s))=\frac{1}{\lambda\mp k_{*}+\sqrt{s_{*}-s}}.$ (42)
Let us now reproduce this result from the integral representation (29).
The key observation here is that as $s$ increases, the branch points at $k=\mathrm{i}A(s)$ and $k=-\mathrm{i}B(s)$ move toward the origin. Once $s=s_{*}-k_{*}^{2}$, we have $B(s)=0$, and the branch point falls precisely on the integration contour. For $s>s_{*}-k_{*}^{2}$ the contour would cross the branch cut, invalidating the integral representation (29). To circumvent this issue, we shift the integration contour from the real axis to a horizontal line $\mathcal{C}_{*}=\left\{k:\,\mathrm{Im}(k)=k_{*}\right\}$ (see Fig. 4). This yields
$\psi^{\pm}_{\text{a.c.}}(\lambda;s)=\exp\left[-\frac{1}{2\pi}\int_{\mathcal{C}_{*}}\mathrm{d}k\,\frac{1}{\lambda\pm\mathrm{i}k}\log\big[(k-\mathrm{i}A)(k+\mathrm{i}B)\big]\right],\quad\,\mathrm{Re}(\lambda)>k_{*}.$ (43)
Note that if $\mathrm{Re}(\lambda)<k_{*}$, then the pole at $k=\mathrm{i}\lambda$ appearing in $\psi^{+}(\lambda;s)$ lies between the original real-axis contour and the deformed contour $\mathcal{C}_{*}$. In this case, the deformation would capture the residue at the pole, introducing an additional factor in (43). We omit these technical details here, as they are not essential for understanding the general procedure, and for simplicity assume that $\mathrm{Re}(\lambda)>k_{*}$. In Section 6, while performing the continuations for $\phi^{\pm}(\lambda;\rho,s)$ in the $s$-plane, we account for this properly.
The representation (43) provides an analytic continuation in $s$ valid up to the critical value $s=s_{*}$, where the branch points coincide, trapping the contour of integration so that no deformation can avoid them. This creates a pinch singularity which in the $s$-plane manifests as a branch point at $s_{*}$. As a side note, such behavior is well-known in the context of analytic structure of the S-matrix in high-energy physics [undefaal].
To verify that (43) coincides with (42), we perform a simple change of variables $k\mapsto k+\mathrm{i}k_{*}$ in (43). This immediately recovers (29) with $A\mapsto A-k_{*}$, $B\mapsto B+k_{*}$ and $\lambda\mapsto\lambda\mp k_{*}$. Computing the integrals as in (39) reproduces (36), thereby confirming the consistency of the contour deformation.
Figure 4. Contour deformation in the $k$-plane. As $s$ increases, the branch point at $k=\mathrm{i}A(s)$ moves downward and the branch point at $k=-\mathrm{i}B(s)$ moves upward. At $s=s_{*}-k_{*}^{2}$, the branch point at $k=-\mathrm{i}B(s)$ falls onto the real line and the integral representation (29) becomes invalid. Shifting the integration contour to $\mathcal{C}_{*}$ at $\mathrm{Im}(k)=k_{*}$ keeps it between the branch points, allowing analytic continuation (43) for $s<s_{*}$.
An overview
Since the technical core of the rest of the paper consists of analytic continuations of the functions $\phi^{\pm}(\lambda;\rho,s)$, we believe it might be helpful to provide a brief overview of the underlying procedure in light of the toy example described above. The continuation strategy is fundamentally the same, but with considerably more technical details. It is thus useful to keep the simplified picture from the toy example in mind while reading Sections 4, 5, 6, and 7 (Section 8 is conceptually different and is not illustrated by this example).
In Section 4, we analyze the analytic structure in the $k$-plane of the integrands defining the functions $\phi^{\pm}(\lambda;\rho,s)$ in (9). We demonstrate that the structure closely resembles that of (29), and the singularities governing the asymptotic behavior are: (i) poles at $k=\mp\mathrm{i}\lambda$ and (ii) two branch cuts. We then perform the analytic continuation in the $\lambda$-plane by tracking the movement of the poles, following the same procedure as in the toy example.
In Section 5, we study the probability $S_{\infty}(X_{0})$ which, according to (4) and (7), is obtained by taking the limit $\rho\to 0$ and $s\to 1$ in (8). In terms of the analytic structure, this limit corresponds to one of the branch points falling onto the real line, as in the toy example at $s=s_{*}-k_{*}^{2}$. This analysis requires careful consideration of the behavior of the integrands in the vicinity of $k=0$.
In Section 6, we perform the continuation in the $s$-plane. Specifically, we show that the asymptotic decay of the survival probability $S_{N}(n\,|\,X_{0})$ is governed by the point $s_{*}$ at which two branch points coincide. This leads to the decay rate (14). In Section 7, we perform a similar procedure for the $\rho$-plane and show that the branch points coincide at $\rho_{*}$, yielding the corresponding decay rate in (14).
Having outlined this strategy, we now proceed to the computations. We encourage the reader to keep this overview and the toy example in mind throughout the technical analysis that follows.